OpenStudy (anonymous):

: 2 radical 3 + 6 radical 27 Pleeaasssee heeellppp!!

OpenStudy (anonymous):

\[2\sqrt{3} + 6\sqrt{27} = 2\sqrt{3}+18\sqrt{3} = 20\sqrt{3}\]

OpenStudy (anonymous):

ok quit!

OpenStudy (anonymous):

Quit what?

OpenStudy (anonymous):

are you using a text editor or can you just type latex in your sleep?

OpenStudy (anonymous):

i meant "ok i quit". not 'ok quit'

OpenStudy (anonymous):

:D I'm the one that needs to retire :D I'm getting a lot wrong for some reason. At least mrinyaya always catches me :D (or watchmath) (or satellite73)